OpenStudy (anonymous):

45 is what percent of 15?

ganeshie8 (ganeshie8):

\(\large \frac{45}{15} \times 100\)

OpenStudy (anonymous):

remember betty when you had 25% of what number is 13? this problem is exactly the same, except now you do it in reverse

OpenStudy (anonymous):

We have : \[\Large\frac{45}{15}=x\%=\frac{x}{100}\] So : \[x=\frac{45}{15}\times100 \\~~=300 \] So 45 is 300% of 15

OpenStudy (anonymous):

so 300 is the answer

OpenStudy (anonymous):

of course !